.5 chance of getting each question right.
4 questions
.5^4= .0625
What is the difference between a probability density curve and cummulative distribution function?
what is density curve
How do you solve a standard normal distribution?
You do not solve a standard normal distribution. It is not a question nor an equation or inequality to be solved. You can answer questions using the standard normal distribution but what you do depends on the question and on what information is given.
Given a Poisson distribution with mean equals 2 Find P x5?
Given a Poisson distribution with mean = 2. Find P(X < 5)
How could the principle of probability be used to predict the results of genetic crosses?
In genetical experiments the principle of probability is used to predict the results at !% and 5% probability levels. If a particular character appears in 95 plants out of 100, it is considered as a significant character to be inherited. For reaching more accuracy it may be tested at 1% level.
One must first consider the number of ordered ways that 4 students may be chosen, and then divide that by the number of ways the four may have been ordered to create a number of distinct groups.
For the orders, there are 30 choices for the first, 29 for the second, 28 for the third, and 27 for the fourth. However, there are a 4 ways within the four to pick the first, 3 ways for the second, 2 for the third, and 1 left for the fourth. The answer may be expressed by the following:
30 x 29 x 28 x 27
______________
1 x 2 x 3 x 4
which evaluates to 27,405 committees.
Gerolamo Cardano (1501 - 1576).
How can you draw a tree diagram to show the results of tossing 4 coins a total of 96 times?
You cannot.
The tree diagram for tossing 4 coins has 16 branches. So if that is done 96 times, you will have a tree with 1696 branches which is approx 4 trillion googol branches.
How many queens of diamonds in a deck of cards?
In an ordinary 52-card deck, there is one Queen of Diamonds.
Suppose you have two random variables, X and Y and their joint probability distribution function is f(x, y) over some appropriate domain.
Then the marginal probability distribution of X, is the integral or sum of f(x, y) calculated over all possible values of Y.
What is the probability of getting a face card or a 7 or a diamond from an ordinary deck of cards?
There is a .53846 probability (53.846%) if aces are counted as face cards. If they are not counted the probability drops to .48077 (48.077%)
P = 1/2 * 1/2 * 1/6 = 1/24
You can get marble from a quarry or a stone merchant, you can also sometimes find marble tiles at a tile and ceramic shop.
What is the binomial theorem formula for probability?
I expect you mean the probability mass function (pmf). Please see the right sidebar in the linked page.
When dealing cards which direction do you deal them?
Try alternating between clockwise and counter-clockwise and see if anyone at the table notices. If someone comments, deal the direction that they want to go to avoid further conflict.
Bob scored 300 or 700.